# Conditional Denoising Diffusion Probabilistic Models for Data Reconstruction Enhancement in Wireless Communications

### v2

@@ II-B Parametric Reverse Diffusion Process
Invoking (9)–(11), one can infer that the covariance matrix in (9) has no learnable parameter. Hence, we simply need to learn the mean vector ${\tilde{\bm{\mu}}}(\bm{x}_{t},\bm{x}_{0},t)$. In the following, we first simplify (10), and then propose the DDPM-based training framework.
Thanks to the reparameterization trick and with a similar approach to (4), we can express ${\bm{x}}_{0}$ as follows
@@
Now we can learn the conditioned probability distribution $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ of the reverse diffusion process by training a neural network that approximates $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Therefore, we simply need to set the approximated mean vector $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ to have the same form as the target mean vector $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Since $\bm{x}_{t}$ is available at time-step $t$, we can reparameterize the neural network to make it approximate $\bm{\epsilon}_{t}$ from the input $\bm{x}_{t}$. Compiling these facts results in the following expression for $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$
@@
Invoking (14) and (15), at each time-step $t$, the DDPM model takes $\bm{x}_{t}$ as input and returns the distortion components $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$. Moreover, $\bm{\epsilon}_{t}$ denotes the diffused noise term at time-step $t$.
@@ III-A System Model & Formulas
where $P$ denotes the transmit power, $\tau^{t}$ is the distortion noise caused by the transmitter hardware with the corresponding impairment level $\kappa^{t}$, and $\tau^{r}$ reflects the hardware distortion at the receiver with $\kappa^{r}$ showing the level of impairment at the receiver hardware. Notably, ${\tau^{r}}$ is conditionally Gaussian, given the channel realization ${h}$. This is a well-known and experimentally-validated model for HIs, which is widely-adopted in wireless communication literature [21]. We emphasize that according to [21], the distortion noise caused at each radio frequency (RF) device is proportional to its signal power, where a fixed portion of the information signal is turned into distortion noise due to quantization errors in analog-to-digital converter (ADC), inter-carrier interference induced by phase noise, leakage from the mirror subcarrier under I/Q imbalance, nonlinearities in power amplifiers, etc. To be aligned with the “batch-processing” nature of AI/ML algorithms and wireless communications simulators [20], we reformulate the signaling expressions in matrix-based format. To do so, we define the batch (with size $K$) of information-bearing samples as ${\bm{s}}\overset{\Delta}{=}\left[s_{1},s_{2},\cdots,s_{K}\right]^{\sf T}$ corresponding to the batch of ground-truth information samples ${\bm{x}_{0}}\overset{\Delta}{=}\left[x_{0}^{(1)},x_{0}^{(2)},\cdots,x_{0}^{(K)}\right]^{\sf T}$ with the underlying distribution $\bm{x}_{0}\sim q(\bm{x}_{0})$. Hence, the batch of received signals at the receiver can be expressed as
@@
To address the abovementioned challenges, we incorporate the noisy degraded signals into the denoising framework of the diffusion model, realizing a conditional DDPM. Accordingly, the denoising process is modified and the corresponding conditional diffusion and reverse processes are derived.
III-B Proposed Solution: Conditional DDPM for Enhanced Data Reconstruction
@@ III-B1 Conditional Forward Process
III-C Conditional Reverse Process
$\mathcal{L}_{t}^{\sf cond}=$
wehre $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$. Intuitively speaking, the loss function in (35) is designed such that the conditional DDPM learns, over the denoising steps, to estimate both the Gaussian noise $\bm{\epsilon}$ and the residual errors of estimating $\widehat{\bm{x}}$, i.e., ${\widehat{\bm{x}}}-{\bm{x}}_{0}$. Therefore, the the coefficients for $({\widehat{\bm{x}}}-{\bm{x}}_{0})$ and $\bm{\epsilon}$ are chosen similar to the coefficient of ${\widehat{\bm{x}}}$ and the standard deviation of the diffusion process as in (27), respectively. According to (29)–(35), the training and sampling algorithms are summarized in Algorithm 1 and Algorithm 2, respectively. A visual description of the training phase is also exhibited in Fig. 3.
@@ III-C Conditional Reverse Process
Figure 3: General description of training the diffusion model.
@@
IV Numerical Results
In this section, we provide different numerical evaluations to highlight the performance of our proposed approach in different scenarios and compared to different baselines. We also evaluate the resilience of our proposed scheme under low-SNR regimes. Our scheme is implemented in conjunction with the NVIDIA Sionna simulator [20] to show the compatibility of the scheme with a practical communication system. To elaborate, after that the diffusion model is trained offline using Algorithm 1, it is plugged into the end of the receiver block of the Sionna simulator, where both scenarios of DNN-based receiver, known as Sionna’s neural receiver, and the conventional receiver are considered in the following evaluations. [^3] One interesting direction would be to propose a framework in which the forward and reverse process of diffusion training are directly incorporated into the communication simulator. This requires modeling the conditional probability distribution of data transmission and reconstruction as modified forward and reverse diffusion processes, respectively, which will be studied in our future works. We consider 64-QAM modulation symbols with 5G low density parity check coding (LDPC) [20] for the modulation and coding of the transmission scheme. We also note that the transmit SNR is defined as
$\Gamma=10\log_{10}(\frac{P}{\sigma^{2}})\hskip 5.69054pt\text{dB}.$ (36)
Without loss of generality, we normalize the transmit power by setting the average signal power to $P=1$ for all experiments, while varying the SNR by setting the standard deviation (std) of noise $\sigma$ [25].
For the choice of neural network for our conditional diffusion model, we follow the model employed in [7], and adopt the U-Net architecture [23] as our denoising neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)$ with input size 32 and depth of 4. The employed neural network is supposed to take the distorted image at each time step, and output the predicted noise with the same size as the input image. The U-Net architecture matches well with this requirement. It consists of a contracting path and an expansive path, which respectively capture contextual information and reduce the spatial resolution of the input, and then upsamples the feature maps, while also performing convolutional operations [23]. Additionally, it introduced residual connections between the encoder and decoder, which greatly improves the gradient flow. Evaluations are carried out on MNIST dataset (resized to $32\times 32$) which contains 60,000 training images and 10,000 testing images of hand-written digits. The corresponding dataset $\mathcal{D}$ in Algorithm 1, is then obtained by transmitting the training samples of MNIST over the Sionna’s simulator and obtain the decoded signals at the receiver’s output, and then pairing each of the ground-truth samples with their degraded version to create $(\bm{x}_{0},{\widehat{\bm{x}}})\in\mathcal{D}$. The training is carried out for 10 epochs over $T=200$ steps, and using adaptive moment estimation (Adam) optimizer with learning rate $\lambda=10^{-3}$. To stabilize the training process, exponential moving average (EMA) method is implemented [24] for model update during training. This helps maintain a form of “model momentum.” Specifically, instead of directly updating model’s weights, a copy of the previous weights is kept, and the weighted average between the previous and new version of the weights are calculated for the model update. We also set $T=200$. For the conditioning weights $10<\lambda_{t}<$ 1, we consider a linear scheduling over $T=200$ steps. For the variance scheduling $\beta_{t}$, we also consider a linear scheduling, starting from $\beta_{0}=0.0001$ to $\beta_{T}=0.0095$. We note that the values for the variance scheduling were obtained based on trial-and-error over different scheduling formats. To the best of our knowledge, there does not exist a unified generic solution for the choices of variance scheduling in the diffusion model literature, and it depends on the specific problem to which the diffusion model is applied. To enable employing only one neural network for the entire denoising steps, instead of training $T$ distinct models for each time-step $t\in[T]$, we share the parameters of the neural network across time-steps. More specifically, as shown in Fig. 3, we encode the time-steps $t\in[T]$ and input it to each hidden layer of our neural network as a positional embedding [7]. The output of the hidden layers are then multiplied by the time embeddings [7, 9, 12]. Intuitively, this makes the model “know” at which particular time-step it is operating for every sample in the batch.
IV-A Data Visualization
@@
Fig. 4 visualizes the performance of our conditional DDPM-aided reconstruction enhancement scheme. For this figure, we consider the extreme condition of $\Gamma=0$ dB channel SNR, with the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. As can be seen from the figure, under such extreme conditions, the naive format of a communication system cannot perform well in removing the noise and distortions to obtain a clean denoised version of the ground-truth information signals (Fig. 4-(b)). In such scenarios, our proposed diffusion-aided scheme can improve the reconstruction of images by removing the noise and distortions conditioned on the degraded signals. This results in a significant improvement in the decoding performance. This improvement is actually obtained via exploiting the prior knowledge learnt by our conditional DDPM. One might argue that in such cases, we can play with the coding rate and employ stronger error correction codes with lower code rates. However, this can significantly reduce the data rate, while stronger error correction codes might not necessarily work well under such extreme conditions either. We note that in the subsequent figures, we numerically study this, and we show that the DDPM outperforms those schemes in low SNR regimes.
@@
Fig. 5 visualizes the output of our proposed diffusion-aided scheme, compared to a DNN-based receiver as benchmark [18]. For this experiment, we set $\Gamma=5$ dB SNR and the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. Notably, compared to the DNN benchmark our scheme highlights a considerably better performance in obtaining clean samples of the information signals, by removing noise and other residual decoding errors. This is due to the fact that the diffusion-based method exploits the generative prior knowledge of the underlying distribution. However, a DNN-based receiver as proposed in [18] does not take the “content”, nor any prior knowledge of the information signals into account when decoding the information bits. We further emphasize that numerical comparisons between our conditional DDPM and the DNN benchmark are provided in the subsequent figures.
IV-B Reconstruction Performance Metrics

### v3

@@ II-A2 Parametric Reverse Diffusion Process
Invoking (9), one can infer that the covariance matrix has no learnable parameter. Hence, we simply need to learn the mean vector ${\tilde{\bm{\mu}}}(\bm{x}_{t},\bm{x}_{0},t)$. Using the reparameterization trick and with a similar approach to (4), we can express ${\bm{x}}_{0}$ as
@@
Now we can learn the conditioned probability distribution $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ of the reverse diffusion process by training a neural network that approximates $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Therefore, we simply need to set the approximated mean vector $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ to have the same mathematical format as the target mean vector $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Hence we have
@@
Invoking (13), at each time-step $t$, the DDPM model takes $\bm{x}_{t}$ as input and returns the distortion components $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$. Moreover, $\bm{\epsilon}_{t}$ denotes the diffused noise term at time-step $t$.
II-B Practical Implications and Potential Impacts on Communication Systems
As discussed above, DDPMs are inherently connected to the task of denoising. This makes them highly relevant to communication systems, where the unwanted noise and interference is a significant challenge. This implies some of the potential impacts of DDPMs on communication systems. Notably, DDPMs as powerful denoising models can be applied to clean noisy signals received over communication channels. The reverse diffusion process could help reconstruct original signals from noisy transmissions, leading to improved signal quality [4]. Further, DDPMs could be integrated into an end-to-end framework, where both the transmitter and receiver are learned jointly. The transmitter could use a forward diffusion process to generate noisy encoding of the transmitted signal, while the receiver could use the reverse process to decode the original data, optimizing the overall communication pipeline for robustness and efficiency [13]. To conclude, we note that theoretical robustness and practical applications of DDPMs, introduced in this section, highlight that they have the potential to play important roles in future communication systems. This can include, but is not limited to, resilient communications, making communication systems more resilient to noise, interference, and distortions, particularly in extreme environments, such as space (non-terrestrial networks), underwater, or urban areas with heavy interference. High-fidelity signal reconstruction is another direct impact of employing DDPMs for communication networks, which can potentially result in clearer voice calls, sharper video transmissions, and more accurate data recovery in practical communication applications.
@@ III-A System Model & Formulas
where $P$ denotes the transmit power, $\tau^{t}$ is the distortion noise caused by the transmitter hardware with the corresponding impairment level $\kappa^{t}$, and $\tau^{r}$ reflects the hardware distortion at the receiver with $\kappa^{r}$ showing the level of impairment at the receiver hardware. Notably, ${\tau^{r}}$ is conditionally Gaussian, given the channel realization ${h}$. This is a well-known and experimentally-validated model for HIs, which is widely-adopted in wireless communication literature [21]. We emphasize that according to [21], the distortion noise caused at each radio frequency (RF) device is proportional to its signal power, where a fixed portion of the information signal is turned into distortion noise due to quantization errors in analog-to-digital converter (ADC), inter-carrier interference induced by phase noise, leakage from the mirror sub-carrier under I/Q imbalance, non-linearities in power amplifiers, etc. To be aligned with the “batch-processing” nature of AI/ML algorithms and wireless communications simulators [22], we reformulate the signaling expressions in matrix-based format. To do so, we define the batch (with size $K$) of information-bearing samples as ${\bm{s}}\overset{\Delta}{=}\left[s_{1},s_{2},\cdots,s_{K}\right]^{\sf T}$ corresponding to the batch of ground-truth information samples ${\bm{x}_{0}}\overset{\Delta}{=}\left[x_{0}^{(1)},x_{0}^{(2)},\cdots,x_{0}^{(K)}\right]^{\sf T}$ with the underlying distribution $\bm{x}_{0}\sim q(\bm{x}_{0})$. Hence, the batch of received signals at the receiver can be expressed as
@@
To address the above-mentioned challenges, we incorporate the noisy degraded signals into the denoising framework of the diffusion model, realizing a conditional DDPM. Accordingly, the denoising process is modified and the corresponding conditional diffusion and reverse processes are derived.
@@
III-B Proposed Solution: CDiff for Enhanced Data Reconstruction
@@ III-B1 Conditional Forward Process
III-B2 Conditional Reverse Process
$\mathcal{L}_{t}^{\sf cond}=$
where $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$. Intuitively speaking, the loss function in (33) is designed such that the conditional DDPM learns, over the denoising steps, to estimate both the Gaussian noise $\bm{\epsilon}$ and the residual errors of estimating $\widehat{\bm{x}}$, i.e., ${\widehat{\bm{x}}}-{\bm{x}}_{0}$. Therefore, the coefficients for $({\widehat{\bm{x}}}-{\bm{x}}_{0})$ and $\bm{\epsilon}$ are chosen similar to the coefficient of ${\widehat{\bm{x}}}$ and the standard deviation of the diffusion process as in (25), respectively. According to (27)–(33), the training and sampling algorithms are summarized in Algorithm 1 and Algorithm 2, respectively. A visual description of the training phase is also exhibited in Fig. 4.
@@ III-B2 Conditional Reverse Process
Figure 4: General description of training the diffusion model and the notion of time embedding. Each time-step is mapped (via positional embedding) to an embedding vector, which is then incorporated into the hidden layers of the neural network. The figure shows time-step $t=21$ as an example.
IV Numerical Results
In this section, we provide different numerical evaluations to highlight the performance of our proposed approach in different scenarios and compared to different baselines. We also evaluate the robustness of our proposed scheme under low-SNR regimes.
IV-A Experimental Setup
IV-A1 Communication Setup
Our scheme is implemented in conjunction with the NVIDIA Sionna simulator [22] to show the compatibility of the scheme with a practical communication system. To elaborate, after that the diffusion model is trained offline using Algorithm 1, it is plugged into the end of the receiver block of the Sionna simulator as shown in Fig. 1. In the evaluations, both scenarios of DNN-based receiver, known as Sionna’s neural receiver, and also the conventional receiver are considered in the evaluations. [^4] One interesting direction would be to propose a framework in which the forward and reverse process of diffusion training are directly incorporated into the communication simulator. This requires modeling the conditional probability distribution of data transmission and reconstruction as modified forward and reverse diffusion processes, respectively, which will be studied in our future works. We consider 64-QAM modulation symbols with 5G low density parity check coding (LDPC) [22] for the modulation and coding of the transmission scheme. The transmit SNR is defined as
$\Gamma=10\log_{10}(\frac{P}{\sigma^{2}})\hskip 5.69054pt\text{dB}.$ (34)
Without loss of generality, we normalize the transmit power by setting the average signal power to $P=1$ for all experiments, while varying the SNR by setting the standard deviation (std) of noise $\sigma$ [24]. Following the experimentally-validated model of [21], we consider the impairment levels in the range $[0,0.15^{2}]$ [4, 1], where smaller values correspond to less-impaired transceiver hardware.
IV-A2 Learning Setup
Evaluations are carried out on MNIST dataset [^5] http://yann.lecun.com/exdb/mnist (resized to $32\times 32$). The dataset is split into 60,000 training images and 10,000 testing images of hand-written digits [25]. The corresponding dataset $\mathcal{D}$ in Algorithm 1, is then obtained by transmitting the training samples of MNIST over the Sionna’s simulator and obtain the decoded signals at the receiver’s output, and then pairing each of the ground-truth samples with their degraded version to create $(\bm{x}_{0},{\widehat{\bm{x}}})\in\mathcal{D}$. For validation of our numerical analysis, mean-squared error (MSE) and PSNR metrics are studied in this paper. These metrics measure the quality of reconstruction, where a higher(lower) PSNR(MSE) indicates better quality.
Training of our scheme can be carried out via two different training strategies, i.e., varying training SNR (simulating the system with different transmit SNRs), or fixed training SNR (fixing the communication scenario to a determined SNR). Our validations in terms of the PSNR metric showed that the performance of the fixed training SNR scenario saturates too early during the validation phase. Therefore, we choose the varying scenario, where we set $-15\leq\Gamma^{\sf train}\leq 15$ dB by trial and error over different ranges of SNRs. Notably, having training samples $(\bm{x}_{0},{\widehat{\bm{x}}})$ from a wide range of communication SNRs and the typical range of transmitter and receiver impairment levels $\kappa^{t,r}\in[0,0.15^{2}]$ (See Table I) indicates that the wireless conditions are implicitly taken into account as the conditioning of our model. This way, we incorporate the conditions of a practical wireless system into account.
For the choice of neural network for our CDiff model, we follow the model employed in [9], and adopt the U-Net architecture [26] as our denoising neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)$ with input size 32 and depth of 4. The employed neural network is supposed to take the distorted image at each time step, and output the predicted noise with the same size as the input image. The U-Net architecture matches well with this requirement. It consists of a contracting path and an expansive path, which respectively capture contextual information and reduce the spatial resolution of the input, and then up-samples the feature maps, while also performing convolutional operations [26]. Additionally, it introduced residual connections between the encoder and decoder, which greatly improves the gradient flow. The training is carried out for 10 epochs over $T=200$ steps, and using adaptive moment estimation (Adam) optimizer with learning rate $\lambda=10^{-3}$. To stabilize the training process, exponential moving average (EMA) method is implemented [27] for model update during training. This helps maintain a form of “model momentum.” Specifically, instead of directly updating model’s weights, a copy of the previous weights is kept, and the weighted average between the previous and new version of the weights are calculated for the model update. For the conditioning weights $10<\lambda_{t}<$ 1, we consider a linear scheduling over $T=200$ steps. For the variance scheduling $\beta_{t}$, we also consider a linear scheduling, starting from $\beta_{0}=0.0001$ to $\beta_{T}=0.0095$. We note that the values for the variance scheduling were obtained based on trial-and-error over different scheduling formats. To the best of our knowledge, there does not exist a unified generic solution for the choices of variance scheduling in the diffusion model literature, and it depends on the specific problem to which the diffusion model is applied. To enable employing only one neural network for the entire denoising steps, instead of training $T$ distinct models for each time-step $t\in[T]$, we share the parameters of the neural network across time-steps. More specifically, as shown in Fig. 4, we encode the time-steps $t\in[T]$ and input it to each hidden layer of our neural network as a positional embedding [9]. The output of the hidden layers are then multiplied by the time embeddings [9, 11, 14]. Intuitively, this makes the model “know” at which particular time-step it is operating for every sample in the batch. To summarize, Tables I and II highlight the experimental setup and hyper-parameters and metrics used for the simulation.
TABLE I: Parameters of the communication setup.
Parameter Value Simulation Setup NVIDIA Sionna Link-level Simulator [22] Modulation scheme 64 QAM [22] Coding scheme 5G LDPC [22] Training SNR $[-15,15]$ dB Impairment level $[0,0.15^{2}]$ [21, 4, 1]
TABLE II: Hyper-parameters of the learning setup.
Parameter Value Dataset MNIST [25] Train-test split 60000 - 10000 Neural architecture U-Net with depth 4 (32, 64, 128, 256) [9, 26] Total denoising steps $T=200$ Training epochs 10 variance scheduling linear for ($\beta_{0},\beta_{T})=(0.001,0.0095)$ [9] Learning rate 0.001 with Adam optimizer [27]
IV-B Baseline methods used for comparison
The following baselines are employed for comparison with our proposed approach.
  • DNN-based receiver [18] incorporated into Sionna’s simulator [^6] https://nvlabs.github.io/sionna/examples/Sionna_tutorial_part2.html. This baseline is studied for both qualitative and quantitative comparisons in Figs. 6 and 8, respectively. For the DNN-based receiver baseline, we are inspired by the network architecture proposed in [18], and since we have employed our diffusion model at the receiver, we only exploit the receiver DNN of [18] and fine-tune it for benchmarking, so that we can have a fair comparison. Three hidden linear layers with $64$ neurons and rectified linear unit (ReLU) activation functions are implemented, trained for 5000 iterations with learning rate $0.01$. Based on our trial and errors, increasing the depth of the conventional DNN, or the number of hidden neurons does not result in any significant improvement in its performance. We show that our proposed model outperforms this baseline since the CDiff exploits the generative prior knowledge regarding the underlying distribution of the ground-truth information signals. However, the DNN receiver does not take the “content”, nor any prior knowledge of the information signals into account when decoding the information bits.
  • VAE-based reconstruction. VAEs as another state-of-the-art GenAI-based decoding approach, which are famous for image reconstruction, are compared with our DDPM approach. This is quantitatively studied in Fig. 7. As mentioned earlier, our approach considers the “contents” of the transmitted signal during denoising and reconstruction. Therefore, to have a fair comparison from the perspective of “prior knowledge” utilization, we consider this baseline, which also considers the prior knowledge about the underlying distribution of the ground-truth when reconstructing. The VAE’s loss function is composed of two parts, i.e., the reconstruction loss and the Kullback-Leibler (KL) divergence loss. The former measures the difference between the original data and the reconstructed output generated by the decoder, while the latter measures the difference between the learned probability distribution and the predefined prior distribution. For the “contracting path,” four 2D convolutional layers (Conv2D) are implemented with $32$, $64$, $64$, and $64$ filters, respectively, with kernel size of $3\times 3$, and ReLU activation functions, followed by a flattening layer and a dense layer of size $32$ (to combine the down-sampled learned features before entering the latent space). Two dense layers further generate the parameters of the Gaussian distribution in the latent space (a.k.a. the mean vector $\mu$ and log variance vector $\sigma$ which represent the compressed latent features of the input), with latent space dimension Dim set to $5$ and $20$ for different scenarios (see Fig. 7). The decoder reconstructs the images from the latent representation, by exactly mirroring the encoder in reverse, by progressively up-scaling the compressed data back to its original dimension. Accordingly, the architecture is implemented by a dense layer and a reshape layer to up-scale to a higher dimension, and then followed by de-convolutional layers (Conv2DTranspose), to further upscale and refine the data to reconstruct the original input. We show the superior performance of our approach compared this baseline, thanks to the high-quality sampling characteristic of diffusion models.
  • Conventional Receiver (Naive scheme). This baseline is studied for qualitative comparison (data visualization in Fig. 5), as well as quantitative comparisons (Figs. 9 and 10). Our proposed scheme complements this baseline in a way that instead of employing error correction codes with low code rates to improve data decoding in extreme channel conditions, our proposed scheme can be employed while preserving the data rate. We show that we can maintain higher data rates while also improving the reconstruction quality. We also show that the functionality of the proposed CDiff can be exploited to compensate for the underlying errors within the noisy data at the receiver, outperforming the channel codes in conventional receivers.
In what follows, we show how our proposed method outperforms or complements the existing approaches.
@@ IV-B Baseline methods used for comparison
IV-C Qualitative Comparisons: Data Visualization
@@
Fig. 5 visualizes the performance of our conditional DDPM-aided reconstruction enhancement scheme. For this figure, we consider the extreme condition of $\Gamma=0$ dB channel SNR, with the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. As can be seen from the figure, under such extreme conditions, the naive format of a communication system cannot perform well in removing the noise and distortions to obtain a clean denoised version of the ground-truth information signals (Fig. 5-(b)). In such scenarios, our proposed diffusion-aided scheme can improve the reconstruction of images by removing the noise and distortions conditioned on the degraded signals. This results in a significant improvement in the decoding performance. This improvement is actually obtained via exploiting the prior knowledge learnt by our conditional DDPM. One might argue that in such cases, we can play with the coding rate and employ stronger error correction codes with lower code rates. However, this can significantly reduce the data rate, while stronger error correction codes might not necessarily work well under such extreme conditions either. In the subsequent figures, we numerically study this and we show that the DDPM outperforms those schemes in low SNR regimes.
@@
Fig. 6 visualizes the output of our proposed diffusion-aided scheme, compared to a DNN-based receiver as benchmark [18]. For this experiment, we set $\Gamma=5$ dB SNR and the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. Notably, compared to the DNN benchmark our scheme highlights a considerably better performance in obtaining clean samples of the information signals, by removing noise and other residual decoding errors. This is due to the fact that the diffusion-based method exploits the generative prior knowledge about the underlying distribution of the ground-truth information signals. However, a DNN-based receiver as proposed in [18] does not take the “content”, nor any prior knowledge of the information signals into account when decoding the information bits. We further emphasize that numerical comparisons between our scheme and the DNN benchmark are provided in the subsequent figures.
IV-D Quantitative Comparisons: Reconstruction Performance